### v1

# LHU-Net: A Light Hybrid U-Net for Cost-Efficient, High-Performance Volumetric Medical Image Segmentation

Abstract
As a result of the rise of Transformer architectures in medical image analysis, specifically in the domain of medical image segmentation, a multitude of hybrid models have been created that merge the advantages of Convolutional Neural Networks (CNNs) and Transformers. These hybrid models have achieved notable success by significantly improving segmentation accuracy. Yet, this progress often comes at the cost of increased model complexity, both in terms of parameters and computational demand. Moreover, many of these models fail to consider the crucial interplay between spatial and channel features, which could further refine and improve segmentation outcomes. To address this, we introduce LHU-Net, a Light Hybrid U-Net architecture optimized for volumetric medical image segmentation. LHU-Net is meticulously designed to prioritize spatial feature analysis in its initial layers before shifting focus to channel-based features in its deeper layers, ensuring a comprehensive feature extraction process. Rigorous evaluation across five benchmark datasets — Synapse, LA, Pancreas, ACDC, and BRaTS 2018 — underscores LHU-Net’s superior performance, showcasing its dual capacity for efficiency and accuracy. Notably, LHU-Net sets new performance benchmarks, such as attaining a Dice score of 92.66 on the ACDC dataset, while simultaneously reducing parameters by $85\%$ and quartering the computational load compared to existing state-of-the-art models. Achieved without any reliance on pre-training, additional data, or model ensemble, LHU-Net’s effectiveness is further evidenced by its state-of-the-art performance across all evaluated datasets, utilizing fewer than 11 million parameters. This achievement highlights that balancing computational efficiency with high accuracy in medical image segmentation is feasible. Our implementation of LHU-Net is freely accessible to the research community on GitHub.
Keywords Volumetric Medical Image Segmentation, Deep Learning, Computational Efficiency, High-Performance Segmentation, Light Hybrid Architecture
1 Introduction
Medical imaging technologies such as MRI, CT, or planar X-ray imaging, have substantially transformed diagnostics by providing non-invasive visualizations of internal anatomical structures. Segmentation of medical imaging data is commonly performed to support diagnostics, intervention planning, or disease progression assessment. However, manual segmentation by radiologists is not only time-consuming, but is also highly subjective and prone to inconsistencies. Accordingly, there is a growing need for automated segmentation approaches.
Recently, deep learning approaches (and Convolutional Neural Networks (CNNs) in particular) have shown great potential for medical image segmentation. Prominent CNN-based segmentation networks include Fully Convolutional Networks (FCNs) Long et al. (2015), U-Net Ronneberger et al. (2015), DeepLab Liang-Chieh et al. (2015), ResNet He et al. (2016), and V-Net Milletari et al. (2016). Each of these network designs has made significant progress by consistently outperforming earlier state-of-the-art networks Azad et al. (2022); Bozorgpour et al. (2023).
Although CNNs can detect local patterns, they are fundamentally unable to comprehend the wider context or global connections present in an image. This constraint becomes especially apparent while performing intricate image analysis tasks that require a comprehensive understanding of the entire scene or image. The Vision Transformer (ViT) Dosovitskiy et al. (2020) design has been central to recent developments aimed at overcoming this constraint. The attention mechanism is the spine of ViT’s architecture which allows the aggregation of data over the whole input sequence. Overcoming the limited receptive field limitations commonly seen in CNNs, this strategy enables the network to absorb long-range contextual information Shamshad et al. (2023). Nevertheless, here a major drawback is set by the underrepresentation of local information hence compromising the performance Azad et al. (2023a). This is more noticeable when it comes to detecting local textures with a fine grain, which is an essential component of many diagnostic and prognostic activities Azad et al. (2023b). This restriction stems from the fact that ViTs use a novel approach to image processing. To comprehend the inter-dependencies among the image’s patches, ViT models use self-attention mechanisms to break the image into smaller pieces. Although ViTs have made advancements in capturing overall contextual information, CNNs may have an advantage in extracting specific features within a local area because of the differences in their fundamental methods for image processing Geirhos et al. (2018).
Consequently, there has been a paradigm shift towards hybrid models that amalgamate the strengths of both CNNs and vision transformer models Geirhos et al. (2018). A notable strategy involves the fusion of CNN and transformer model features, leveraging the local feature extraction capabilities of CNNs and the global contextual understanding of transformers. This synergy aims to overcome the individual limitations of each architecture. For instance, TransUNet Chen et al. (2021), an innovative architecture, merges transformer layers within the CNN bottleneck, enabling the simultaneous modeling of both local and global dependencies. Another exemplary model, HiFormer Heidari et al. (2023), introduces the integration of a Swin transformer module with a CNN-based encoder. This architecture generates dual multi-scale feature representations, which are then coalesced through a Double-Level Fusion module, enhancing the overall segmentation accuracy. Further, architectures like UNETR Hatamizadeh et al. (2022) and CoTr Xie et al. (2021) have been instrumental in advancing 3D medical image segmentation. UNETR utilizes a transformer-based encoder in tandem with a CNN decoder, optimizing the segmentation process. CoTr, on the other hand, innovatively bridges a CNN encoder and decoder with a transformer, specifically targeting an enhanced performance in low-resolution stages of segmentation. Furthermore, TransBTS Wang et al. (2021) represents another stride in this direction, employing a similar hybrid approach to optimize segmentation results. Additionally, architectures such as Swin-UNet Cao et al. (2022) offer unique advantages. The Swin transformer’s shifted window-based self-attention mechanism within a U-Net-like architecture facilitates the efficient handling of large-scale medical images while capturing intricate details and broad structural contexts, making it highly effective for complex segmentation tasks.
CNNs are distinguished by their hierarchical architecture, adept at discerning spatial hierarchies in imagery via convolutional filters, benefiting from their inherent translation invariance and localized connectivity. This design tends to yield models characterized by a reduced parameter count due to the sharing of weights across spatial domains and the confinement of convolution operations to localized areas. As a result, CNNs generally demand less computational power, particularly in applications where spatial coherence is pivotal. Conversely, ViTs employ self-attention mechanisms, which, although offering comprehensive analysis capabilities across an image, cause a quadratic surge in computational costs relative to the input size. This attribute renders ViTs more demanding in terms of computationally and parameter volume. Hybrid models seek to amalgamate CNNs’ local processing efficiencies with ViTs’ expansive global receptive fields. Yet, this ambition often culminates in a marked escalation of parameters and computational expenditure. Certain hybrid models indiscriminately increase model complexity, adding layers or expanding attention mechanisms across the network in the hope of boosting performance. This brute-force scaling-up strategy often results in diminishing returns on performance gains relative to the higher computational overhead Zhou et al. (2021); Yang et al. (2024); Chen et al. (2023); Shaker et al. (2022). The insufficiency of current hybrid models, particularly in the context of volumetric medical image segmentation, may well stem from their sub-optimal self-attention configurations. Moreover, these models typically fail to explicitly leverage the interplay between spatial and channel features, an oversight that could otherwise improve segmentation accuracy. A refined strategy adapts attention mechanisms to suit various model stages. Initially, where the acquisition of spatial information is crucial, specialized attention blocks tailored for spatial focus can effectively manage local features. As the model delves deeper, the emphasis shifts towards assimilating broader contextual data, necessitating attention frameworks that prioritize channel dynamics. Such sequential customization facilitates a harmonious balance between model intricacy and efficacy. Aligning architectural design with the specific demands of information processing at each stage enables enhanced performance, circumventing the disproportionate parameter and computational load increase seen in more homogenized hybrid models. In essence, the evolution of hybrid models offers a promising avenue to exploit the unique strengths of both CNNs and ViTs. Nevertheless, simply amplifying model complexity fails to serve as a panacea. Instead, a meticulously crafted network featuring bespoke attention mechanisms attuned to the distinct requirements of varying processing phases emerges as a more adept and resource-efficient methodology for capitalizing on the synergistic potential of CNNs and ViTs in visual data analysis.
In this vein, we introduce LHU-Net (Light Hybrid U-Net), a model embodying the ideal equilibrium between complexity and performance. LHU-Net exemplifies that achieving an optimal balance is not only feasible but pivotal, setting a new precedent in the design of hybrid models for sophisticated image segmentation tasks.
Figure 1: Accuracy (Dice score) vs. model complexity (FLOPs and parameters) comparison on ACDC.
Contributions
This work introduces several significant advancements in the realm of medical image segmentation, pushing the boundaries of current methodologies and setting new benchmarks for efficiency and accuracy. Below, we delineate our key contributions, which underscore the potential of our proposed LHU-Net in addressing complex segmentation tasks across diverse medical imaging datasets.
Efficient Hybrid Attention Selection for Better Contextual Understanding: A cornerstone of our approach is the strategic deployment of two specialized attention mechanisms within ViTs, meticulously engineered to calibrate the model’s focus across varying stages of analysis. Incorporating Large Kernel Attention with an additional deformable layer (LKAd) in conjunction with spatial and channel attention, this deliberate synthesis empowers our model to adeptly navigate long-range dependencies. This approach not only facilitates the extraction of expansive global contextual insights but also addresses texture bias, ensuring the preservation of high-frequency details without sacrificing the integrity of feature representation and inductive biases. In the initial phases, our architecture employs spatial attention to masterfully discern local features and subtleties within the early layers of the network. As the model evolves, it seamlessly transitions to channel attention, a pivotal shift that underscores the importance of distilling and accentuating critical features within the deeper layers. This bifurcated attention strategy enables LHU-Net to harness the strengths of both spatial and channel attentions effectively, facilitating a nuanced and comprehensive feature extraction that is attuned to the unique demands of medical image segmentation.
Setting New Efficiency Benchmarks in Medical Segmentation: LHU-Net represents a paradigm shift in achieving high-performance segmentation with significantly reduced computational resources. In one notable dataset (Automated Cardiac Diagnosis Challenge), our architecture achieved a reduction in the number of parameters by approximately 85% and a quartering of FLOPS (see Figure˜1), all while maintaining a lean parameter footprint of under 11 million across all five evaluated datasets. This breakthrough demonstrates that a well-crafted, purpose-specific architecture can achieve exceptional performance without the traditional trade-offs between computational demand and accuracy. It proves that balancing model complexity with computational efficiency, far from being an elusive goal, is a tangible reality with LHU-Net.
Versatile Superiority Across Diverse Imaging Modalities and Segmentation Challenges: LHU-Net’s design principles and execution excellence have propelled it to the forefront of state-of-the-art methodologies across multiple datasets, showcasing unparalleled versatility. Whether analyzing CT scans, MR images, or datasets that incorporate up to four distinct modalities (as in the BraTS challenge), LHU-Net consistently delivers top-tier performance. Furthermore, its adaptability extends to handling a wide range of segmentation tasks, from single-label segmentation to complex scenarios involving up to eight distinct labels (e.g., in the MICCAI Multi-Atlas Abdomen Labeling Challenge). This broad spectrum applicability underscores LHU-Net’s robustness and its potential as a universal solution for medical image segmentation, capable of accommodating the multifaceted nature of clinical imaging data.
In summary, our contributions through the development of LHU-Net not only advance the technical capabilities in medical image segmentation but also offer a scalable, efficient pathway to integrating deep learning solutions into clinical workflows. By addressing the critical challenges of computational efficiency and model adaptability, LHU-Net paves the way for more accessible and effective diagnostic tools, promising significant impacts on patient care and medical research.
Figure 2: The proposed LHU-Net architecture. On the left, an abstract representation of the proposed network is depicted. At the top levels, the ‘Init’ and ‘Out’ blocks manage input and output respectively, while ‘CNN Macro Blocks’ facilitate information flow between these top-level blocks and the ‘Hybrid Macro Blocks’ at the network’s core. Thorough explanations of the utilized attention blocks and their sub-components, along with other relevant details, are provided on the right side of the network, offering a comprehensive understanding of the architecture’s design. Note that ‘PE’ and ‘RB’ stand for Positional Embedding and ResBlock, respectively.
2 Methodology
Our approach, named LHU-Net, is designed to efficiently extract enriched features across various network levels while optimizing parameter efficiency. By integrating convolutional-based blocks with hybrid attention mechanisms, LHU-Net effectively captures contextual information, addressing both strong local features and non-local dependencies. Figure˜2 illustrates the architecture of LHU-Net, showcasing its hierarchical U-Net encoder-decoder structure. The encoder stages, encode 3D input patches into varying levels of representations, spanning from local to global, facilitating the transmission of essential information via skip connections to the decoder stages, ultimately generating the final segmentation masks. Table˜1 shows the notations used throughout this section.
Table 1: Notations used throughout Section˜2.
Acronym Meaning
BN Batch normalization layer
ViT Vision transformer
PW-Conv Point-wise convolution block
DW-Conv Depth-wise convolution block
DWD-Conv Depth-wise dilated convolution
DDW-Conv Deformable depth-wise convolution
GELU Gaussian error linear unit
LKAd Large kernel attention with one additional deformable layer
A Self-attention matrix
n Number of tokens ($HWD$)
h The number of heads in the ViT attention block
x Input variable
$d$ Hidden state dimension
Q The query embedding inputs
K The key embedding inputs
V The value embedding inputs
S The skip connection
F The feature map
ResBlock Residual Block
PReLU Parametric rectified linear unit
PE Positional embedding
CA Channel attention module
SA Self-attention module
Network Architecture: LHU-Net comprises three distinct stages respecting the spatial dimensions of the feature map. The first stage includes the “Init Block,” and “Out Block,” at the beginning and the end of the network consists of a single block maintaining the spatial dimensions of both the input and output of the network. This stage plays a pivotal role in preserves the spatial shape of the input tensor throughout the network, ensuring the accurate transformation of input data for the encoder and producing the desired output after the decoder. At the intermediate stage, convolutional blocks (termed “CNN Blocks”) are utilized to reduce spatial dimensions while enhancing local features, capturing high-frequency representations effectively. Subsequently, the output is channeled to the final stage (named “Hybrid Blocks”), where a trio of hybrid attention mechanisms, incorporating convolutional large kernel attention Guo et al. (2023) with one additional deformable layer Azad et al. (2024) alongside spatial and channel-wise ViT attention blocks, further refines features, accounting for global representations and long-range dependencies. Throughout the encoder stages, skip connections maintain connectivity with corresponding decoder blocks, ensuring seamless information flow. Detailed explanations of each stage follow in subsequent subsections.
2.1 Init and Out Blocks
To optimize the performance of subsequent convolutional blocks in the deeper levels of the network and enhance the overall quality of the final output, special emphasis is placed on the boundary regions by employing the Init and Out stages. These stages ensure that all processing is conducted within the same spatial space as the input image, preserving its spatial dimensions while boosting the model’s effectiveness. The Init block begins with a point-wise convolutional operation (PW-Conv) applied to the input data, adjusting the channel dimensions to match those of the subsequent level.
$F_{\text{I}}=\text{BN}(\text{PReLU}(\text{PW-Conv}(I)))$ (1)
Here, $F_{\text{I}}$ is of the same spatial size as $I$ (the input image). It represents the feature map produced directly from the input image, serving as the input for subsequent levels and acting as the skip connection for the last step of the CNN Decoder Block simultaneously.
Moreover, the model applies a ResBlock directly to the input data, facilitating the extraction of high-frequency representations that encapsulate sensitive local features. This approach enables the capture of crucial details for accurate boundary estimation in the final segmentation output. Subsequently, these representations are fed into the Out block (Equation˜2), contributing to the establishment of optimal guidance for the final mask ($X_{\text{Out}}$).
$X_{\text{Out}}=\text{PW-Conv}(\text{ResBlock}(\text{ResBlock}(I)+X_{\text{CD}}))$ (2)
$X_{\text{CD}}$ denotes the output of the CNN Decoder Block.
2.2 CNN Blocks
An integral part of our proposed model are pure CNN blocks, renowned for their adeptness in capturing robust local contextual features at these high- and mid-frequency feature levels. Unlike fully ViT approaches, our method strategically integrates ResBlocks in two levels to diminish spatial dimensions without overwhelming feature representation, preparing the input for the subsequent hybrid stage. This decision is based on two considerations: Firstly, the initial spatial dimensions carry substantial computational costs and don’t efficiently leverage spatial ViT blocks, while also failing to achieve the comprehensive channel representation needed for effective channel-wise ViT attention blocks. Secondly, by refraining from downsampling in the Init block, our focus remains on optimizing parameter efficiency and preserving local features, delegating the responsibility for more global features to the subsequent level.
At each level of encoders ($l$), the Down Conv block encodes the input feature $F_{l}$ into the output feature $F_{l+1}$ as defined:
$\begin{aligned} \hat{F}&=\text{BN}(\text{DW-Conv}(\text{LeakyReLU}(\text{BN}(\text{DW-Conv}_{\text{Down}}(F_{l})))))\\ R&=\text{BN}(\text{PW-Conv}(F_{l}))\\ F_{l+1}&=\text{LeakyReLU}(\hat{F}+R)\end{aligned},$ (3)
where, $\text{DW-Conv}_{\text{Down}}$ denotes the convolution operation responsible for downsampling the spatial dimensions of $_{l}. F^\hat{F}$ represents the processed feature map obtained by applying normalization, convolution, and activation functions to the downsampled feature map $F_{l}$. $R$ captures the original input feature $F_{l}$ after a normalization and convolutional operation. Finally, $F_{l+1}$ is computed by adding the processed feature map $\hat{F}$ to the original feature map $R$, followed by an activation function. Also, Equation˜3 can be considered A ResBlock operations when $R=F_{l}$. This process facilitates the flow of information while efficiently managing spatial dimensions across different levels of the encoder.
2.3 Hybrid Blocks (Hybrid Attention)
The strategic incorporation of the LKAd module within the network architecture, particularly in intermediate to deep layers, is instrumental in enhancing the network’s efficacy for semantic segmentation tasks, wherein the intricate interplay between local details and global context is paramount for accurate delineation of semantic boundaries and objects within the image.
In this final stage of the proposed network, comprising intermediate and deep layers, the emphasis shifts towards employing hybrid attention techniques. Having established a robust foundation in capturing and learning local features in the preceding stage, our focus now broadens to encompass a diverse array of attention mechanisms.
This includes the integration of larger kernel sizes, which enable to capture a broader range of spatial representations, coupled with deformations to ensure the preservation and appropriate focusing on the desired receptive fields. Simultaneously, incorporating spatial ViT attention in the initial layers of this stage and channel ViT attention in the final layers serves to aggregate rich and high-level descriptions within the channels, facilitating the extraction and incorporation of optimal features and descriptions.
This strategic arrangement not only provides efficiency by judiciously selecting appropriate combinations across different layers, but also avoids the need for repeated use of a fixed block Shaker et al. (2022); Rahman et al. (2024); Liu et al. (2024) or multiple repetitions Azad et al. (2024); Wang et al. (2023a); Hong et al. (2023); Li et al. (2023) across levels. Such an approach makes it possible to capture nuanced and meaningful descriptions while maintaining complexity and minimizing the number of parameters, thereby facilitating effective learning.
2.3.1 Self-Adaptive Contextual Fusion Module
Incorporating the spatial attention block within the top hybrid blocks, both residually in the encoder stage and collectively in the decoder stage, alongside the convolutional information flow, presents a sophisticated strategy to fortify the model’s capacity in addressing global structures. This integration facilitates the effective capture of global contextual information through a fusion of spatial attention maps, computed in parallel using the outputs of LKAd and self-attention mechanisms. By strategically intertwining these components, this block aids in mitigating information loss, especially within the deeper layers of the encoder, while seamlessly transferring features to the CNN Blocks in the decoder.
In the spatial attention block, we leverage the LKAd block alongside spatial self-attention. This strategic combination enables the model to concurrently address long-range dependencies for extracting comprehensive global contextual representations and mitigate texture bias while ensuring the effective extraction of features and inductive biases without compromising the high-frequency representation. LKAd facilitates the consideration of large regions within the input image in a flexible and adaptive conventional manner, thereby enabling a broader receptive field by combining the information with a deformable grid which is utilized after two convolution layers of LKAd compared to standard attention mechanisms confined to fixed-size regions. Such adaptability empowers the block to efficiently capture both local and global contextual information.
As depicted in Figure˜2, the final output of the Self-adaptive Contextual Fusion block is computed as:
$F_{S}=\text{Comb}\left(F,\delta_{s}\mathbf{LKAd}(F)+\gamma_{s}\mathbf{SA}(F)\right)$ (4)
Where $F\in\mathbb{R}^{C\times H\times W\times D}$ represents the input feature maps, and $F_{S}$ denotes the output of this block with the same dimensions. $\delta_{s}$ and $\gamma_{s}$ represent learnable parameters per channel that control the combination weights of the two different attention mechanisms.
LKAd: This module orchestrates attention computation, as delineated in Azad et al. (2024) by the following formulation:
$\hat{x}$ $=\text{GELU}(\text{PW-Conv}(x))$ (5)
[…]
$x_{\text{LKAd}}$ $=\text{PW-Conv}(\text{DDW-Conv}_{3}(x_{\text{LA}}))$
$x_{\text{Ad}}$ $=\text{PW-Conv}(x_{\text{LKAd}}\otimes\hat{x})+x,$
where the input tensor $x\in\mathbb{R}^{C\times H\times W\times D}$ undergoes a series of transformations aimed at enhancing its representational capacity and contextual understanding. First, the input is subject to a pointwise convolution operation, denoted as $\text{Conv}_{1}$, followed by an activation function, Gaussian Error Linear Units (GELU), to induce nonlinearity and dimensionality reduction. The transformed tensor $\hat{x}$ is then processed through a sequence of depth-wise convolutions (DW-Conv) and depth-wise dilated convolutions (DWD-Conv) to extract multi-scale features while preserving spatial information. The resulting tensor $x_{\text{LA}}$ encapsulates locally attentive representations. Notably, the distinctive aspect of the LKAd module lies in its integration of deformable depth-wise convolutions (DDW-Conv), denoted as $\text{DDW-Conv}_{3}$, which introduces deformable grids to enable adaptive sampling of feature maps, thereby enhancing the model’s ability to capture fine-grained details and long-range dependencies. Finally, acknowledging that $\otimes$ and ++ denote element-wise multiplication and summation, respectively, the refined feature map $x_{\text{LKAd}}$ is convolved with the original input $\hat{x}$ through a pointwise convolution operation, facilitating the fusion of global contextual information with local features. This resulting tensor, $x_{\text{Ad}}$, embodies a synergistic fusion of local and global context, effectively enriching the representation of the input tensor $x$.
Self-Attention: The $SA$ refers to the self-attention module. To compute the final result $SA(x)$, the normalized tensor $x$ of shape $n,C$ is generated, where $n=HWD$. This involves computing $Q$, $K$, and $V$ projections using 3 linear layers, resulting in dimensions $n,h,C/h$, where $h$ denotes the number of heads, by performing $Q=W^{Q}X$, $K=W^{K}X$, and $V=W^{V}X$. Here, $W^{Q}$, $W^{K}$, and $W^{V}$ represent the projection weights for $Q$, $K$, and $V$, respectively. To manage complexity and learn spatial information efficiently, a technique proposed in Shaker et al. (2022) is used, which involves two projections for $K$ and $V$. This yields $K_{p}=W^{p}_{K}\times K$ and $V_{p}=W^{p}_{V}\times V$, projecting the $n$ quantized spatial tokens from $K$ and $V$ into $p$ learnable complex spatial separations in $K_{p}$ and $V_{p}$, altering the dimensions from $n,h,C/h$ to $p,h,C/h$, where $p\ll n$. Then, by multiplying the normalized $Q$ by the transpose of $K_{p}$ and calibrating the resulting values of each head using the learnable parameter $\alpha$, and applying a softmax function ($\sigma$), the similarity between each feature and the remaining spatial features is obtained. Finally, these computed similarities are multiplied by the projected $V$ layer to produce the final spatial attention with dimensions $n\times C$. This spatial attention mechanism is described by the equation:
$x_{SA}=\sigma\left(\frac{QK_{p}^{T}}{\sqrt{d}}\right)\cdot V_{p}$ (6)
Here, $Q$, $K_{p}$, and $V_{p}$ stand for queries, projected keys, and projected spatial value layers, respectively, each of size $d$. Next, we pass $x_{SA}$ through a linear normalization followed by a linear layer to enhance the performance as the result of $SA$ module. Finally, through the integration of the Comb function, as defined in Equation 7:
$\text{Comb}(x,A)$ $=C+\text{PW-Conv}(\text{DW-Conv}_{3}(C))$ (7)
$\text{where}\quad C$ $=x\otimes(1+A)$
where $\text{DW-Conv}_{3}$ is the 3×3×3 convolution block, we obtain the final output.
2.3.2 OmniFocus Attention Block
Located at the deepest level of our network architecture (which has the highest number of feature layers of all hybrid blocks), both residually in the encoder stage and collectively in the decoder stage, this block synergizes with the convolutional information flow to provide a crucial enhancement in feature representation to augment the model’s ability to capture intricate channel-wise relationships and helps the model gain the ability to selectively emphasize relevant meaningful features while suppressing noise across spatial locations. The integration of ViT channel attention along with LKAd in this omni focus attention block, enables the efficient capture of interdependencies between different channels through the use of the proposed combined attention mechanisms, facilitating seamless feature transfer for the next decoder layers. Within this attention block, we leverage the LKAd block, similar to Self-Adaptive Contextual Fusion Module, alongside channel attention mechanisms. This synergistic combination enables the model to simultaneously address inter-channel relationships and long-range dependencies, facilitating the extraction of comprehensive feature representations. By integrating these mechanisms, the channel attention block empowers the model to effectively extract and utilize rich contextual information for enhanced feature representation and improved final performance.
As depicted in Figure˜2, the final output of this attention block similar to Equation˜8 is computed as:
$F_{S}=\text{Comb}\left(F,\delta_{c}\mathbf{LKAd}(F)+\gamma_{c}\mathbf{CA}(F)\right)$ (8)
Where $F$ denotes the previously extracted input feature maps in the shape of $C,H,W,D$ and $\theta_{c}$ and $\gamma_{c}$ represent learnable parameters per channel layers that control the combination weights of the two different attention mechanisms with respect to each head ($h$). The LKAd module computes attention to the input feature $F$ following the formulation introduced in Equation˜5, while $CA$ denotes the channel attention module. This module captures interdependencies between feature channels by employing the dot product operation in the channel dimension between the channel value layer and the channel attention maps. Similar to the creation of $Q$, $K$, and $V$ projections using three linear layers for self-attention, resulting in dimensions $n,h,C/h$, we repeat this process to obtain $Q_{C}$, $K_{C}$, and $V_{C}$ for calculating channel attention:
$x_{CA}=V_{C}\cdot\sigma\left(\frac{Q_{C}^{T}K_{C}}{\sqrt{d}}\right)$ (9)
Next, we pass $x_{CA}$ through a linear normalization followed by a linear layer to improve performance, akin to the result of the $CA$ module. Then, as defined in Equation˜7, using the given input $x$ and the summation of weighted attention results, we generate the final output utilizing the Comb function.
3 Experimental setup and analysis
In this section, we discuss datasets, experimental setup, and evaluation matrices, used to showcase the effectiveness of our proposed framework for medical image segmentation.
3.1 Datasets
MICCAI Multi-Atlas Abdomen Labeling Challenge (Synapse) Landman et al. (2015): The Synapse dataset consists of 30 abdominal CT scans ranging in size from 512 x 512 x 85 to 512 x 512 x 198 voxels. This collection offers a wide range of slice thicknesses (2.5 mm to 5.0 mm) and in-plane resolutions (0.54 x 0.54 mm² to 0.98 x 0.98 mm²), making it ideal for developing and testing advanced 3D segmentation algorithms. Following the method described in Azad et al. (2024); Shaker et al. (2022), the training set of the dataset consists of 18 samples, with the remaining 12 samples reserved for testing. Each scan in this dataset is precisely labeled to identify eight major abdominal organs: the aorta, gallbladder, spleen, left and right kidneys, liver, pancreas, and stomach. The images were randomly cropped to 128 × 128 × 64 for model input, as described in Zhou et al. (2021); Isensee et al. (2021). The same method as in previous studies Zhou et al. (2021); Chen et al. (2023) was used for inference.
Automated Cardiac Diagnosis Challenge (ACDC) Bernard et al. (2018): The ACDC dataset, specifically designed to advance 3D cardiac image segmentation, consists of a collection of cine MRI images from a diverse group of 100 patients. This dataset is crucial for the advancement of advanced 3D segmentation algorithms, specifically in the accurate delineation of three major cardiac anatomical structures: the left ventricle (LV), the right ventricle (RV), and the myocardium (Myo). Each scan in this dataset provides intricate visual representations of these structures, allowing for extremely precise segmentation. The MRI scans vary in spatial resolution, typically ranging from 1.37 to 1.68 mm² per pixel. Furthermore, the dataset consists primarily of images with a slice thickness of 5 mm, while there are some occurrences with a slice thickness of 8 mm thickness. Occasionally, there is a 5 mm gap between slices. The variance in the thickness and the gaps between slices present specific obstacles for segmentation algorithms. These challenges require the use of strong approaches that can successfully handle these differences.
In terms of preprocessing, the method adheres to the protocols established in Chen et al. (2023); Shaker et al. (2022), involving the random cropping of images to dimensions of 160 x 160 x 16 and using sliding windows for the inference. The patient dataset is partitioned into separate subsets for the purpose of model training and evaluation. Specifically, 70 patients are assigned for training, 10 for validation, and 20 for testing. This allocation ensures a thorough evaluation of the segmentation models that are generated using the ACDC dataset.
Left Atrial Dataset (LA) Xiong et al. (2021): constitutes a curated set of 100 three-dimensional MR images, all of which have undergone gadolinium enhancement. These images are uniform in resolution, each measuring 0.625 mm across all three spatial dimensions, with two prevalent spatial configurations: 640 × 640 × 88 and 576 × 576 × 88 pixels. The dataset has been divided into five distinct groups, with each group consisting of 20 patients. During the preprocessing stage, we followed the preprocessing in Luo et al. (2021); Wang et al. (2023b) where the images are standardized to a zero mean and unit variance to facilitate consistent model input. The images also undergo strategic cropping, incorporating augmented margins to ensure the focal region is accentuated. During model training, random cropping of the volumes is performed to yield sub-volumes of 96 × 96 × 96, which serve as inputs for the model. During the inference step, The study followed the methods outlined by Wang et al. (2023b) utilizing the sliding window technique provided by the MONAI frameworkCardoso et al. (2022).
NIH pancreas dataset (CT-82) Roth et al. (2015): This dataset encompasses a collection of 82 publicly available, contrast-enhanced, annotated three-dimensional CT scans of the abdominal region. Each scan within this dataset is characterized by a resolution of 512 × 512 pixels across the axial plane, with a variable number of slices spanning from 181 to 466 to accommodate the diverse anatomical structures present in the patient population. For structured analysis, the dataset is organized into four cohorts, with each segment comprising 20, 20, 21, and 21 patients, respectively. In alignment with the preprocessing methods proposed by Luo et al. (2021); Wang et al. (2023b), this dataset underwent a standardized preprocessing protocol. The same configurations for training and inference as established for the LA dataset were applied.
Brain Tumor Segmentation (BraTS) Challenge 2018 Bakas et al. (2017); Menze et al. (2014): The Brain Tumor Segmentation (BraTS) Challenge datasets from the year 2018 play a pivotal role in advancing medical image analysis. This dataset, constituting a diverse array of pre-operative Magnetic Resonance Imaging (MRI) scans, are sourced from a multitude of institutions. Key characteristics of these datasets include the rigorous alignment of the MRIs, isotropic resampling to a resolution of 1 × 1 × 1 mm, and the removal of the skull (skull-stripping). The MRIs present a uniform size of 240 × 240 × 155. The data, encompassing a broad spectrum, are collected from several academic centers using varied MRI scanning equipment. Each dataset encompasses four MRI modalities: native T1, post-contrast T1-weighted (T1c), T2-weighted (T2), and T2 Fluid-attenuated Inversion Recovery (FLAIR). Tumor segmentation in these datasets is categorized into three distinct sub-regions: the Whole Tumor (WT), the Tumor Core (TC), and the Enhancing Tumor (ET). Prior to model training and analysis, this dataset underwent a standardized preprocessing method, following the preprocessing guideline proposed by Dorent et al. (2019). For the training phase, MRI scans were randomly cropped to a size of 128 × 128 × 128 voxels to fit the network input. During the inference phase, we employed MONAI’s Cardoso et al. (2022) sliding window technique.
The BraTS-2018 dataset includes MRIs from 285 patients, with 210 exhibiting high-grade gliomas and 75 displaying low-grade gliomas. The data allocation was as follows: 70% for training, 10% for validation, and 20% for testing. The 2018 BraTS datasets provide an unparalleled resource for the advancement of automated brain tumor segmentation techniques, presenting diverse challenges and opportunities for the development of cutting-edge image analysis methods.
3.2 Experimental setup
The developed neural network model is implemented using Python (version 3.10.12) and is built upon the PyTorch framework Paszke et al. (2019); Falcon and The PyTorch Lightning team (2019) (version 2.1.0). The model was trained on a single NVIDIA A100 graphics processing unit, equipped with 80 GB of VRAM. The training leveraged a composite loss function, DICE loss Milletari et al. (2016) with binary cross-entropy, expressed as:
$\mathcal{L}_{total}=0.5\cdot\mathcal{L}_{dice}+0.5\cdot\mathcal{L}_{ce}$ (10)
In terms of optimization, we have utilized the Stochastic Gradient Descent (SGD) optimization method. The SGD configuration includes a weight decay factor of $3\times 10^{-5}$, and a momentum of 0.99, enhanced with Nesterov momentum and a varied learning rate for each dataset. The learning rate is dynamically updated at each epoch following the equation:
$\text{learning rate}=\text{base }lr\times\left(\frac{epoch}{max\_epoch}\right)^{0.9}$ (11)
Table 2: Configuration specifics for network training across various datasets, highlighting the tailored approach to optimize performance for each unique dataset challenge.
Synapse ACDC BraTS 2018 LA & Pancreas
Crop size 128 x 128 x 64 160 x 160 x 16 128 x 128 x 128 96 x 96 x 96
Base learning rate 0.003 0.003 0.01 0.01
Batch size 2 4 2 2
Sliding window overlap (%) 50 50 50 85
Downsample [^1] A single integer indicates uniform downsampling across all axes. [2, 2, 1], 2, 2, 2, 2 [2, 2, 1], [2, 2, 1], [2, 2, 1], 2, 2 2, 2, 2, 2, 2 2, 2, 2, 2, 2
Data augmentation techniques consistent with Zhou et al. (2021); Shaker et al. (2022) were employed. The training duration varied across datasets, with the Synapse and ACDC datasets undergoing 1000 epochs, featuring 250 patches per epoch, amounting to 250,000 iterations in total. Conversely, for the BraTS datasets, a uniform 1000 epochs were implemented, whereas the LA and Pancreas datasets were subjected to 10,000 iterations, reflecting the tailored approach to training regimen based on dataset specificity. Table˜2 encapsulates the network configurations deployed for each dataset under study, detailing parameters such as crop size, initial learning rate, batch size, percentage of sliding window overlap, and downsampling strategy. These configurations underscore the customized approach to each dataset, optimizing for both performance and computational efficiency.
3.3 Evaluation metrics
Dice Similarity Coefficient (DSC) Milletari et al. (2016): It is used to quantify the similarity between the predicted segmentation and the ground truth in medical images, formulated as below:
$\text{DSC}=\frac{2\times|\text{GT}\cap\text{PR}|}{|\text{GT}|+|\text{PR}|}$ (12)
where GT represents the ground truth segmentation, PR is the predicted segmentation, and $|\text{GT}\cap\text{PR}|$ is the number of pixels or voxels common to both.
Jaccard Similarity Coefficient Taha and Hanbury (2015): This coefficient, also referred to as Intersection over Union, measures the overlap between the predicted segmentation and ground truth. It is calculated as:
$\text{Jaccard Index}=\frac{|\text{GT}\cap\text{PR}|}{|\text{GT}\cup\text{PR}|}$ (13)
where $|\text{GT}\cup\text{PR}|$ represents the total number of unique pixels or voxels in either the ground truth or the prediction.
95% Hausdorff Distance (HD95) Taha and Hanbury (2015): It is a metric employed to assess the discrepancy between two point sets, commonly representing the ground truth and the predicted segmentation in image segmentation tasks. Contrary to the traditional Hausdorff Distance, which identifies the maximum distance among the sets, HD95 concentrates on the 95th percentile of these distances to minimize the effect of outliers. Formally, HD95 is delineated as the maximum of two directional distances at the 95th percentile:
$\text{HD95}(\text{GT},\text{PR})=\max\left(\tilde{h}_{95}(\text{GT},\text{PR}),\tilde{h}_{95}(\text{PR},\text{GT})\right)$ (14)
where $\tilde{h}_{95}(\text{GT},\text{PR})$ signifies the 95th percentile of all distances from points in the ground truth set to their nearest counterparts in the prediction set, and $\tilde{h}_{95}(\text{PR},\text{GT})$ indicates the 95th percentile of distances for the inverse relation, from points in PR to their closest points in GT. metric focuses on the closest 95% of points between predicted and actual segmentations, excluding the most extreme 5% to reduce outlier impact. This provides a reliable gauge of segmentation accuracy, reflecting true model performance.
Average Surface Distance (ASD) Taha and Hanbury (2015): ASD measures the average distance from points on the predicted segmentation surface to the closest points on the ground truth surface, and vice versa. The formula is:
$\text{ASD}(\text{GT},\text{PR})=\frac{1}{|S_{\text{GT}}|+|S_{\text{PR}}|}\times\left(\sum_{g\in S_{\text{GT}}}\min_{p\in S_{\text{PR}}}d(g,p)+\sum_{p\in S_{\text{PR}}}\min_{g\in S_{\text{GT}}}d(p,g)\right)$ (15)
where $S_{\text{GT}}$ and $S_{\text{PR}}$ are the surfaces of the ground truth and prediction, respectively, and $d(g,p)$ is the Euclidean distance between points $g$ and $p$.
4 Results and Discussion
4.1 Quantitative Results and Discussion
Table 3: Comparative analysis of our approach versus state-of-the-art in the benchmarks for MICCAI Multi-Atlas Abdomen Labeling Challenge (Synapse). The highest-performing result is highlighted in blue, while the second-best result is denoted in red. Parameter counts are reported in millions (M), and FLOPS are expressed in billions (G). The Dice Similarity Coefficient (DSC) for Spleen (Spl), Right Kidney (RKid), Left Kidney (LKid), Gallbladder (Gal), Liver (Liv), Stomach (Sto), Aorta (Aor), and Pancreas (Pan) are detailed as percentages. Furthermore, the average DSC and 95% Hausdorff Distance (HD95) for all organs are included to provide a comprehensive evaluation of segmentation accuracy. Two-dimensional (2D) models are indicated with *, and three-dimensional (3D) models are italicized.
[…]
DSC $\uparrow$ HD95 $\downarrow$
U-Net* Ronneberger et al. (2015) - - 86.67 68.60 77.77 69.72 93.43 75.58 89.07 53.98 76.85 -
TransUNet* Chen et al. (2021) 96.07 M 88.91 G 85.08 77.02 81.87 63.16 94.08 75.62 87.23 55.86 77.49 31.69
SSTrans-Net* Fu et al. (2024) 29.64 M - 92.23 83.33 86.99 72.87 94.61 81.37 86.72 64.98 82.89 15.55
FET* Azad et al. (2023c) 47.01 M - 91.85 80.06 86.66 69.67 94.43 80.34 85.31 67.08 81.92 18.41
HiFormer* Heidari et al. (2023) 25.51 M - 90.99 79.77 85.23 65.69 94.61 81.08 86.21 59.52 80.39 14.70
UNETR Hatamizadeh et al. (2022) 92.49 M 75.76 G 85.00 84.52 85.6 56.3 94.57 70.46 89.8 60.47 78.35 18.59
Swin-UNETR Hatamizadeh et al. (2021) 62.83 M 384.2 G 95.37 86.26 86.99 66.54 95.72 77.01 91.12 68.8 83.48 10.55
nnFormer Zhou et al. (2021) 150.5 M 213.4 G 90.51 86.25 86.57 70.17 96.84 86.83 92.04 83.35 86.57 10.63
UNETR++ Shaker et al. (2022) 42.96 M 47.98 G 95.77 87.18 87.54 71.25 96.42 86.01 92.52 81.1 87.22 7.53
MoAFormer Luo et al. (2022) $\sim$ 50 M - 92.21 85.23 86.72 71.84 95.92 81.56 89.73 75.46 84.80 16.3
TC-CoNet Chen et al. (2023) 313.67 M 699.58 G 91.77 87.92 88.16 62.00 96.35 79.4 92.45 72.78 83.86 9.64
MS-TCNet Ao et al. (2024) $\sim$ 59 M $\sim$ 172 G 93.70 86.66 86.88 66.96 95.82 82.82 91.60 68.19 84.08 8.44
MOSformer Huang et al. (2024) $\sim$ 77 M $\sim$ 100 G 92.29 83.58 90.32 71.90 95.96 87.87 88.95 74.14 85.63 13.4
HCA-former Yang et al. (2024) $\sim$ 100 M - 92.46 86.67 87.21 70.19 96.52 85.91 92.65 82.24 86.92 7.86
LHU-Net 10.52 M 51.75 G 96.01 87.80 87.57 72.43 96.68 85.79 92.42 81.21 87.49 4.15
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ground Truth LHU-Net nnFormer UNETR++ UNETR
Figure 3: Qualitative comparison of the results of our model with respect to other state-of-the-art (SOTA) in MICCAI Multi-Atlas Abdomen Labeling Challenge (Synapse) (images were generated using 3D Slicer Fedorov et al. (2012)).
MICCAI Multi-Atlas Abdomen Labeling Challenge (Synapse): Table˜3 presents a comprehensive comparison of the LHU-Net model against leading models in the field on the Synapse dataset, a part of the MICCAI Multi-Atlas Abdomen Labeling Challenge. LHU-Net distinguishes itself with an impressive overall average DSC of 87.49%, showcasing its superior organ segmentation accuracy. Remarkably, LHU-Net achieves this high level of accuracy while maintaining a model size of only 10.52 million parameters and requiring 51.74 billion FLOPS, demonstrating an exceptional balance between performance and computational efficiency. Specifically, LHU-Net leads in spleen segmentation with a DSC of 96.01%, evidencing its capability in accurately delineating this organ. It also achieves notable success in segmenting the right kidney, with DSCs of 87.80% illustrating its adeptness in handling organs with complex anatomical structures. The model also performs commendably in segmenting the gallbladder and liver, organs known for their challenging variability in shape and size, achieving DSCs of 72.42% and 96.68%, respectively. This is particularly significant given the gallbladder’s propensity for variable positioning and the liver’s large spatial occupancy, further emphasizing LHU-Net’s robust feature extraction capabilities. A pivotal aspect of LHU-Net’s success is its achievement of the lowest HD95 across all models at 4.15 mm. This demonstrates the model’s enhanced precision, ensuring that 95% of the points on the predicted segmentation surface are within 4.15 mm of the actual segmentation surface. Such precision is critical in clinical contexts, where the accuracy of segmentation has immediate implications for treatment planning and patient outcomes. Contrasted with other high-performance models, LHU-Net not only surpasses them in segmentation accuracy but also in computational efficiency. It outperforms more computationally intensive models like TC-CoNet and Swin-UNETR, demonstrating that increased model complexity does not necessarily correlate with better performance.
Table 4: Comparative analysis of our approach versus state-of-the-art in the benchmarks for Automated Cardiac Diagnosis Challenge (ACDC). The highest-performing result is highlighted in blue, while the second-best result is denoted in red. Parameter counts are reported in millions (M), and FLOPS are expressed in billions (G). Dice Similarity Coefficient (DSC) is provided for the Left Ventricle (LV), Right Ventricle (RV), and Myocardium (Myo), expressed as a percentage. Two-dimensional (2D) models are indicated with *, and three-dimensional (3D) models are italicized. The overall average DSC across all heart regions is also provided to reflect comprehensive model performance.
[…]
TransUNet* Chen et al. (2021) - - 95.73 84.54 88.86 89.71
Swin-UNet* Cao et al. (2022) - - 95.83 85.62 88.55 90.00
SSTrans-Net* Fu et al. (2024) 29.64 M - 94.40 87.51 89.02 90.31
UNETR Hatamizadeh et al. (2022) 92.69 M 77.84 G 94.02 86.52 85.29 88.61
nnFormer Zhou et al. (2021) 37.16 M 47.73 G 95.59 89.53 90.22 91.78
MoAFormer Luo et al. (2022) 46.12 M - 95.15 88.49 90.15 91.26
TC-CoNet Chen et al. (2023) 86.79 M 145.16 G 95.47 88.98 90.27 91.58
D-LKA Net Azad et al. (2024) 39.95 M 42.41 G 94.68 88.04 91.35 91.36
neU-Net Yang et al. (2023) 161.28 M 129.83 G 95.66 89.91 90.75 92.11
HCA-Former Yang et al. (2024) $\sim$ 100 M - 95.8 89.76 90.77 92.11
MOSformer Huang et al. (2024) 77.09 M 100.06 G 96.05 89.65 90.86 92.19
MS-TCNet Ao et al. (2024) 59.49 M 172.89 G 95.77 89.09 89.43 91.43
D-Former Wu et al. (2023) $\sim$ 40 M $\sim$ 50 G 95.93 89.60 91.33 92.29
LHU-Net 5.07 M 12.37 G 96.26 90.56 91.15 92.66
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Image Ground Truth LHU-Net nnFormer TC-CoNet UNETR
Figure 4: Qualitative comparison of the results of our model with respect to other SOTA in Automated Cardiac Diagnosis Challenge (ACDC) dataset.
Automated Cardiac Diagnosis Challenge (ACDC): Table˜4 provides a detailed comparative analysis of advanced deep learning models applied to the Automated Cardiac Diagnosis Challenge (ACDC) dataset, focusing on the segmentation of key cardiac structures: the left ventricle (LV), myocardium (Myo), and right ventricle (RV). Utilizing the DSC as the primary metric for assessment, the LHU-Net model distinguishes itself by attaining an average DSC of 92.66%. This exemplary performance underscores its efficacy in cardiac structure segmentation. The efficiency of LHU-Net is underscored by its lean parameter count of only 5.07 million and the lowest computational footprint, denoted by 12.37 billion FLOPS. This efficiency starkly contrasts with other contenders, where an uptick in parameters and computational complexity has yielded only incremental performance enhancements. LHU-Net’s architecture proves that through thoughtful design, significant leaps in performance can be achieved without proportionally escalating computational demands. LHU-Net showcases superior performance in Myocardium (Myo) segmentation with a DSC of 90.56%, a pivotal achievement considering the Myo’s intricate structure and crucial role in diagnostic evaluations. Its accomplishments in LV and RV segmentation are equally noteworthy, with DSCs of 96.26% and 91.15%, respectively, either surpassing or closely rivalling the top models in these domains. Comparatively, LHU-Net’s nearest rivals, such as D-Former and MOSformer, report impressive DSC scores but incur substantial parameter and computational costs. D-Former, for instance, achieves an average DSC of 92.29% yet demands significantly more resources than LHU-Net, highlighting the diminishing performance returns as model complexity increases. This makes LHU-Net a preferable alternative for medical image segmentation, balancing efficiency with efficacy. The overall percentage improvement in DSC by LHU-Net, alongside a remarkable reduction in parameters by over 85% and FLOPS by almost 70% compared to certain models, marks a notable leap forward in segmentation technology. This efficiency does not detract from accuracy but rather demonstrates that judicious architectural innovations can result in superior performance outcomes.
Table 5: Comparative analysis of our approach versus state-of-the-art in the benchmarks for Left Atrial Dataset (LA). The top-performing result is highlighted in blue, while the next best is marked in red. A five-fold cross-validation approach has been employed, and the results are displayed as mean followed by standard deviation across all folds. Parameter counts are reported in millions (M), and FLOPS are expressed in billions (G). DSC and Jaccard metrics are reported in percentages.
[…]
ASD $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ Jaccard $\uparrow$
SwinUNETR Hatamizadeh et al. (2021) 62.2 M 319.38 G 1.90 (0.34) 91.20 (1.07) 6.07 (1.16) 83.92 (1.57)
SwinUNETR-V2 He et al. (2023) 72.77 M 345.23 G 1.74 (0.28) 91.48 (0.79) 5.59 (1.03) 84.41 (1.26)
UNETR Hatamizadeh et al. (2022) 92.78 M 73.51 G 3.60 (0.47) 88.25 (1.31) 11.55 (1.58) 79.32 (1.97)
UNETR++ Shaker et al. (2022) 29.54 M 29.74 G 1.77 (0.30) 91.05 (1.12) 5.47 (1.01) 83.62 (1.79)
nnFormer Zhou et al. (2021) 149.25 M 176.45 G 2.49 (0.36) 88.79 (1.98) 7.09 (0.75) 80.21 (2.72)
VNet Milletari et al. (2016) 9.44 M 82.76 G 1.99 (0.31) 89.90 (1.79) 6.11 (1.09) 81.91 (2.65)
LHU-Net 8.55 M 32.55 G 1.55 (0.14) 91.55 (0.95) 4.97 (0.75) 84.54 (1.51)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ground Truth LHU-Net nnFormer SwinUNETR-V2 UNETR UNETR++
Figure 5: Qualitative comparison of the results of our model with respect to other SOTA in Left Atrial Dataset (images were generated using 3D Slicer Fedorov et al. (2012)).
Left Atrial Dataset (LA): Table˜5 presents a comprehensive quantitative analysis across several methods. The LHU-Net model establishes itself as a benchmark of excellence in the comprehensive analysis of the Left Atrial Dataset. This model not only excels in standard metrics like the DSC and Jaccard index but also sets new standards in ASD and HD95, providing a holistic view of its segmentation prowess. LHU-Net, with its lean architecture requiring merely 8.57 million parameters and 34.28 billion FLOPS, emerges as the epitome of efficiency and efficacy. This is highlighted by its outstanding average DSC of 91.55, which marginally outperforms SwinUNETR-V2. This underscores LHU-Net’s superior ability to accurately delineate the complex anatomy of the left atrium. This slight yet significant edge in DSC is complemented by leading performance in the Jaccard index at 84.54, reinforcing the model’s effectiveness in capturing the true extent of atrial structures. Crucially, LHU-Net’s excellence extends to the ASD and HD95 metrics, where it achieves the lowest scores of 1.55 and 4.97, respectively. These figures represent the model’s unparalleled precision in approximating the true surface of the atrium and minimizing the distance to the farthest outliers, essential for clinical applications where every millimeter could be critical. The ASD score, in particular, showcases LHU-Net’s ability to maintain consistent proximity to the actual atrial boundary across all dataset variations. Comparatively, while other models like UNETR++ and SwinUNETR-V2 exhibit commendable performances, particularly in 3D segmentation tasks, they do not achieve the same balance of high accuracy and computational efficiency as LHU-Net. For instance, despite SwinUNETR-V2’s competitive DSC and Jaccard scores, LHU-Net’s superior ASD and HD95 metrics underscore its ability to provide more reliable and precise segmentations, which are crucial for clinical diagnostics and interventions.
NIH pancreas dataset (CT-82): Table˜6 presents the quantitative assessment of various segmentation models on the NIH Pancreas dataset (CT-82). LHU-Net’s proficiency is underscored not only by its exemplary performance in the average DSC and HD95 but also in its achievements across the Jaccard index and ASD, metrics essential for assessing the comprehensive accuracy and proximity of predicted segmentations to ground truth. With a minimal architecture demanding only 8.57 million parameters and 34.28 billion FLOPS, LHU-Net achieves remarkable efficiency, evidenced by its top-ranking average DSC of 81.96 and the lowest HD95 at 6.78. This indicates an unparalleled ability to delineate pancreatic tissues accurately, minimizing outliers and extreme segmentation errors that are critically important in clinical diagnostics. The model’s excellence extends to the Jaccard index, achieving a leading score of 70.12, which reflects a high degree of overlap between the segmented and actual pancreatic regions. This metric, along with the model’s superior ASD score of 1.31, highlights LHU-Net’s ability to closely approximate the surface of the pancreas, ensuring that the segmented boundary closely matches the true anatomical boundary, a testament to the model’s precision in capturing intricate organ shapes. In comparison to its counterparts, such as UNETR++ and SwinUNETR-V2, LHU-Net not only presents a more computationally efficient framework but also ensures that this efficiency does not come at the cost of segmentation accuracy. Its balanced approach in optimizing both computational resources and segmentation fidelity positions LHU-Net as an ideal solution for clinical deployment, where time and computational efficiency are as valued as accuracy and reliability. In essence, LHU-Net demonstrates high accuracy across DSC, HD95, Jaccard, and ASD metrics, along with computational efficiency.
Table 6: Comparative analysis of our approach versus state-of-the-art in the benchmarks for NIH pancreas dataset (CT-82). The top-performing result is highlighted in blue, while the next best is marked in red. A four-fold cross-validation approach has been employed, and the results are displayed as mean followed by standard deviation across all folds. Parameter counts are reported in millions (M), and FLOPS are expressed in billions (G). DSC and Jaccard metrics are reported in percentages.
[…]
ASD $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ Jaccard $\uparrow$
SwinUNETR Hatamizadeh et al. (2021) 62.2 M 319.38 G 1.99 (0.51) 80.52 (2.13) 7.52 (1.52) 68.17 (2.79)
SwinUNETR-V2 He et al. (2023) 72.77 M 345.23 G 2.42 (0.97) 81.05 (1.90) 8.85 (2.55) 68.87 (2.52)
UNETR Hatamizadeh et al. (2022) 92.78 M 73.51 G 4.25 (0.83) 76.18 (2.55) 12.98 (2.50) 62.51 (3.24)
UNETR++ Shaker et al. (2022) 29.54 M 29.74 G 1.95 (0.22) 81.49 (0.84) 7.64 (2.01) 69.07 (1.16)
nnFormer Zhou et al. (2021) 149.25 M 176.45 G 3.28 (0.90) 78.05 (2.29) 10.65 (3.74) 64.81 (2.81)
VNet Milletari et al. (2016) 9.44 M 82.76 G 2.79 (0.30) 79.31 (1.95) 8.99 (1.96) 66.42 (2.55)
LHU-Net 8.55 M 32.55 G 1.31 (0.14) 81.96 (1.00) 6.78 (0.52) 70.12 (1.29)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ground Truth LHU-Net nnFormer SwinUNETR-V2 UNETR UNETR++
Figure 6: Qualitative comparison of the results of our model with respect to other SOTA in NIH pancreas dataset (images were generated using 3D Slicer Fedorov et al. (2012)).
Brain Tumor Segmentation (BraTS) Challenge 2018: Table˜7 provides a detailed comparative analysis of various deep learning models, highlighting their performance in terms of DSC and HD95 for Brain Tumor Segmentation (BraTS) Challenge 2018 dataset. In this comparative study, the LHU-Net model stands out as the most proficient, achieving the highest average DSC and the lowest average HD95. This underscores its effectiveness in accurately identifying tumor boundaries and overall tumor regions while minimizing the distance between the predicted and actual tumor borders. LHU-Net’s superior performance in ET and WT segmentation, as well as its exceptional ability to delineate the TC, is further highlighted by its leading scores in these categories. Notably, LHU-Net accomplishes this with a significantly lower parameter count of 10.5 million and a reasonable computational demand, as indicated by its FLOPS measurement. The LHU-Net model’s performance in segmenting the TC is particularly impressive, outperforming all other models with a DSC of 87.42. This is pivotal considering the TC’s clinical importance, as it includes the bulk of the tumor’s malignant tissue. In the WT category, which represents the entire tumor area, LHU-Net again outshines the competition with a DSC of 89.57, indicating its exceptional ability to capture the full extent of the tumor’s presence. In the realm of computational efficiency, LHU-Net demonstrates a commendable balance, achieving high accuracy with relatively low computational overhead. While the UNETR++ model exhibits competitive performance, particularly in the ET segmentation, LHU-Net’s balanced excellence across all measures of tumor segmentation signifies a robust model that can provide comprehensive and accurate segmentations.
4.2 Qualitative Results and Discussion
Following the quantitative analysis, we shift our focus to a qualitative evaluation of the same datasets. Here, we assess the performance of models through visual examination, considering the nuances that are not captured by numerical data alone. This section offers a narrative interpretation of how the models’ segmentations align with the expected outcomes and the ground truth of the medical images. Through this qualitative lens, we glean insights into the models’ practical capabilities, which further informs their potential for real-world clinical integration.
Table 7: Comparative analysis of our approach versus state-of-the-art in the benchmarks for Brain Tumor Segmentation (BraTS) Challenge 2018. The top-performing result is highlighted in blue, while the next best is marked in red. Parameter counts are reported in millions (M), and FLOPS are expressed in billions (G). The terms "Whole Tumor" (WT) refer to the entire tumor area, "Tumor Core" (TC) denotes the central region of the tumor, and "Enhancing Tumor" (ET) highlights areas with active tumor cells that are visible after contrast enhancement.
Methods Params FLOPS DSC $\uparrow$ HD95 $\downarrow$ Average
ET TC WT ET TC WT DSC $\uparrow$ HD95 $\downarrow$
ResUNet Lee et al. (2017) 28.44 M 472.62 G 78.6 82.74 88.75 6.519 9.886 9.547 83.36 8.651
ResUNetSE Toubal et al. (2020) 29.31 M 472.8 G 77.84 82.55 88.67 6.699 8.954 8.066 83.02 7.906
3D-UNET Çiçek et al. (2016) 16.47 M 516.71 G 78.02 82.99 88.92 7.103 9.884 8.228 83.31 8.405
SwinUNETR-V2 He et al. (2023) 72.77 M 836.26 G 80,75 84.15 88.04 5.591 7.807 7.84 84.31 7.079
SwinUNETR Hatamizadeh et al. (2021) 62.2 M 774.98 G 79.43 82.39 88.51 5.835 7.625 10.87 83.44 8.11
UNETR Hatamizadeh et al. (2022) 102.45 M 184.76 G 76.97 78.81 86.51 9.182 11.48 13.44 80.76 11.367
UNETR++ Shaker et al. (2022) 42.65 M 73.59 G 82.29 85.87 89.38 4.348 6.456 6.23 85.85 5.678
nnFormer Zhou et al. (2021) 37.75 M 106.45 G 78.95 85.57 89.4 5.676 6.362 8.723 84.64 6.920
LHU-Net 10.50 M 81.96 G 81.16 87.42 89.57 3.779 5.632 6.481 86.05 5.597
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Image Ground Truth LHU-Net nnFormer SwinUNETR-V2 UNETR UNETR++
Figure 7: Qualitative comparison of the results of our model with respect to other SOTA in Brain Tumor Segmentation (BraTS) Challenge 2018 dataset.
MICCAI Multi-Atlas Abdomen Labeling Challenge (Synapse): Figure˜3 offers a qualitative perspective on the segmentation capabilities of various models in the MICCAI Multi-Atlas Abdomen Labeling Challenge (Synapse). For the first patient, visualized in the top row, we observe that UNETR struggles significantly, producing fragmented and scattered segmentations that fail to represent the organs as cohesive entities. nnFormer and UNETR++ exhibit irregular segmentation contours, particularly for the left kidney and aorta, while UNETR++ additionally misses the stomach entirely. In contrast, the LHU-Net model demonstrates a commendable segmentation performance, accurately delineating most organs with smooth boundaries, albeit slightly overestimating the extent of the aorta. The middle row, showcasing the second patient, presents a more challenging scenario where all models encounter difficulties, particularly with the segmentation of the kidneys. Here, both nnFormer and UNETR++ erroneously detect large areas far from the ground truth, highlighting the clinical significance of minimizing false positives. In this complex case, LHU-Net outperforms other models by correctly segmenting other organs and only minor inaccuracies in kidney localization. This indicates a relatively better understanding of the organ boundaries within the abdominal cavity. For the third patient, depicted in the bottom row, the pancreas poses significant segmentation challenges, further complicated by the absence of the gallbladder in the ground truth segmentation. LHU-Net adeptly delineates the organ boundaries precisely, whereas UNETR++ and nnFormer encounter difficulties in accurately segmenting the tail of the pancreas and sharply defining its edges. UNETR, consistent with its performance in previous cases, struggles with segmentation accuracy across multiple organs, including the aorta. Additionally, all models erroneously predict the presence of the gallbladder, highlighting a common difficulty in correctly identifying missing organs across the evaluated models. This observation underscores the necessity for enhanced model sensitivity to the variability and complexities inherent in medical image segmentation tasks. In conclusion, the LHU-Net model exhibits superior performance in organ detection and segmentation across the Synapse dataset. Its segmentation results are not only smoother and more accurate but also showcase robustness against false positives, a critical factor in clinical applicability.
Automated Cardiac Diagnosis Challenge (ACDC): Figure˜4 presents a qualitative, two-dimensional evaluation of cardiac segmentation across several state-of-the-art models within the Automated Cardiac Diagnosis Challenge (ACDC) dataset. This visual assessment spans across three distinct patient cases, each represented by a different row of images, to illuminate the comparative performance of each model in segmenting the left ventricle (LV), myocardium (Myo), and right ventricle (RV). In the first patient, depicted in the top row, there is a noticeable challenge for the models in segmenting the RV, with LHU-Net demonstrating a partial success in capturing some aspects of this anatomical structure. TC-CoNet and UNETR, in particular, exhibit difficulties in accurately delineating the Myo and LV, as evidenced by the deviations from the ground truth. Moving to the second patient, shown in the middle row, the absence of the RV in the slice poses a test of the models’ specificity, where false positives are a concern. In this instance, LHU-Net shows commendable restraint, avoiding the overprediction seen in nnFormer and TC-CoNet, which falsely identify a substantial RV presence. Meanwhile, all models perform well on the Myo and LV, indicating their effectiveness in these regions under less challenging circumstances. The third patient’s imagery, the bottom row, provides a scenario where LHU-Net’s segmentation closely aligns with the ground truth, effectively capturing all relevant cardiac structures. Conversely, nnFormer and TC-CoNet tend towards over-segmentation of the RV, and UNETR presents with under-segmentation, highlighting a potential limitation in their segmentation algorithms. In conclusion, the visualization suggests that LHU-Net possesses superior robustness and precision in cardiac structure segmentation, outperforming the other models under various conditions. This robustness is particularly notable in instances of potential false positives or challenging RV delineation, where LHU-Net maintains high fidelity to the ground truth.
Left Atrial Dataset (LA): Figure˜5 presents a qualitative, three-dimensional visualization of various models’ performance in segmenting the Left Atrial dataset for individual patients. Each row in the figure is dedicated to the segmentation results from a single patient’s dataset, visually assessing each model’s predictive capabilities. In the case of the first patient, shown in the top row, all models except UNETR manage to generate a single entity segmentation, highlighting their ability to integrate disparate anatomical features into a unified model prediction. However, a common challenge across all models is the accurate segmentation of the right corner of the atrium, suggesting a shared area of potential improvement. For the second patient, delineated in the second row, it is discernible that competing models struggle to accurately segment the entirety of the ground truth, resulting in multiple segmented instances rather than a single cohesive structure. Contrarily, LHU-Net succeeds in rendering a single, contiguous segmentation that closely resembles the ground truth, demonstrating its ability to produce smoother and more coherent segmentations. The last patient, as illustrated in the bottom row, presents a challenging case where all models exhibit a tendency for over-segmentation at the base of the atrium. This is indicative of the intrinsic complexity of accurately delineating this anatomical region. Nevertheless, it is worth noting that certain models, specifically UNETR and SwinUNETR-V2, additionally exhibit under-segmentation in other areas, further complicating the segmentation task. In summation, the visualizations in Figure˜5 suggest that LHU-Net consistently generates more robust and smoother segmentations, avoiding the creation of extraneous entities.
NIH pancreas dataset (CT-82): Figure˜6 provides a three-dimensional visualization for the qualitative analysis of pancreatic segmentation across several state-of-the-art models. Each row visually represents the segmentation results from an individual patient’s dataset, offering insight into the predictive precision of each model. For the first patient, illustrated in the first row, it is observable that all models, except for LHU-Net, UNETR++, and SwinUNETR-v2, could not produce a singular, cohesive entity representing the organ. UNETR++ generates a single entity; however, it significantly under-segments the pancreas, failing to capture the complete anatomical structure. In stark contrast, LHU-Net achieves a segmentation that closely approximates the ground truth, underscoring its capability to preserve the organ’s integrity in the segmentation process. The second patient’s data, depicted in the second row, further accentuates the proficiency of LHU-Net in producing a smooth and robust segmentation, particularly when compared to its counterparts. Notably, UNETR++ exhibits a fragmentation in the segmentation, resulting in multiple entities and suggesting a lack of robustness within this dataset. The visualization of the third patient, in the third row, reveals the intrinsic challenges posed by this dataset, as all models predict multiple entities. Nonetheless, LHU-Net distinguishes itself by accurately segmenting the predominant portion that aligns with the ground truth and generating only a minor secondary entity. This contrasts with the other models, which produce multiple, larger erroneous segments, thus compromising the segmentation’s fidelity. Overall, LHU-Net demonstrates superior robustness and smoothness in its segmentation outputs, consistently predicting the pancreas with a higher degree of anatomical accuracy.
Brain Tumor Segmentation (BraTS) Challenge 2018: Figure˜7 provides a qualitative comparison of segmentation results on the Brain Tumor Segmentation (BraTS) Challenge 2018 dataset, illustrating the performance of the LHU-Net model relative to other contemporary state-of-the-art models. Each row in the figure corresponds to a different patient’s MRI slice, showcasing the segmentation of three distinct tumor regions: the necrotic and non-enhancing tumor core (NCR), the peritumoral edematous/invaded tissue (ED), and the enhancing tumor (ET). These regions are critical for evaluating the extent and aggressiveness of brain tumors. In the analysis of the first patient, represented in the first row, a challenge emerges for the models to differentiate between NCR and ET accurately. While many models exhibit proficiency in segmenting the whole tumor (WT) and tumor core (TC), discerning the ET from the NCR proves more problematic. LHU-Net, although striving for accuracy, reveals difficulties at the boundaries of the NCR. SwinUNETR-V2 uniquely attempts to delineate the NCR at the edges, but misplaced detections highlight the complexity of this task. Nevertheless, LHU-Net presents more robust and smoother segmentations overall, indicating its advanced capability for nuanced differentiation within the tumor regions. The second patient, depicted in the second row, presents a scenario with a large NCR and a smaller ET region encircling it. LHU-Net distinguishes itself by accurately delineating all regions, whereas models like UNETR struggle with the finer details of the NCR and ET. Other models tend to overestimate the NCR size, demonstrating the difficulty in capturing precise tumor boundaries. LHU-Net’s performance in this context underscores its exceptional capability to segment with precision, even in complex scenarios. For the third patient, shown in the bottom row, the small NCR and ET amidst the irregularly shaped ED tissue present a significant segmentation challenge. Most models, including UNETR, produce a continuous segmentation of the ED, yet UNETR also segments an additional entity far from the tumor’s actual location, indicating a false positive detection. In contrast, UNETR++ and LHU-Net show a tendency to over-segment the NCR, while SwinUNETR-V2 and nnFormer strive to provide complete segmentation across all regions, despite the challenging nature of the task.
5 Ablation studies
This section is dedicated to dissecting the architectural nuances of our LHU-Net model, specifically focusing on the variations of macro CNN and hybrid modules, as well as the individual contributions of distinct attention mechanisms. These ablation studies are vital to understanding the impact of each architectural component on the model’s performance.
Influence of macro module composition: The LHU-Net model’s architecture employs a blend of convolutional neural networks (CNN) and hybrid macro modules, sequenced strategically for optimal feature extraction. We investigate the model’s performance sensitivity to the depth of the network by varying the number of downsampling stages and the composition of macro modules at each level. In Table˜8, we present a comprehensive analysis of the influence that varying the number of macro modules exerts on the DSC within the ACDC dataset. As established in Section˜2, our configuration preserves the sequence of initially deploying CNN modules followed by hybrid modules. Our findings indicate that reducing the number of downsampling layers from five to four significantly decreases the parameter count, but it adversely affects the DSC, suggesting a trade-off between model complexity and segmentation accuracy. Moreover, we observe that merely increasing the number of CNN macro modules does not enhance performance, confirming that an excess of such modules does not equate to better results. The optimal configuration for our LHU-Net, which is also the proposed model configuration, comprises two CNN macro modules succeeded by three hybrid macro modules. This particular arrangement is instrumental in achieving the model’s pinnacle DSC, thus highlighting the importance of a balanced macro-module composition for sophisticated segmentation tasks.
Table 8: Ablation analysis on the effects of macro module composition on the ACDC dataset performance, evaluated using the average DSC. The optimal model configuration is highlighted in bold.
N. CNNs N. hybrids Params Avg. DSC $\uparrow$
2 2 3.37 M 92.43
2 3 5.07 M 92.65
3 2 3.07 M 92.31
Influence of attention mechanisms within hybrid macro modules: In the hybrid macro modules of LHU-Net, we have the opportunity to deploy different forms of attention blocks—namely, OmniFocus attention and self-adaptive contextual fusion module. OmniFocus attention mechanisms focus on amplifying informative features across channels, while the self-adaptive contextual fusion module prioritizes relevant spatial information. The strategic placement of these blocks within the hybrid macro modules is crucial to the model’s performance. Our study evaluates the efficacy of various configurations of these attention blocks through their impact on the average DSC on the ACDC dataset. Table˜9 presents the results of this analysis, with the optimal sequence of hybrid blocks being two self-adaptive contextual fusion modules followed by an OmniFocus attention block. This configuration yielded the highest average DSC, indicating a superior segmentation capability when spatial attention is prioritized in the initial stages, potentially due to its effectiveness in capturing the spatial distribution of salient features. The subsequent channel attention seems to complement this by refining the feature maps across the channels, thus solidifying the importance of a heterogeneous but strategic arrangement of attention mechanisms.
Table 9: Ablation analysis investigating the impact of attention block sequence within hybrid macro modules on segmentation performance in the ACDC dataset. S represents self-adaptive contextual fusion module, and C represents OmniFocus attention. The most effective sequence is highlighted in bold.
Sequence of attentions Avg. DSC $\uparrow$
SSS 92.32
CCC 92.24
SSC 92.65
SCC 92.5
The evidence from our ablation study demonstrates that the configuration of attention blocks within the hybrid modules is not merely a matter of preference but rather a substantive decision that significantly influences the model’s performance.
6 Conclusion
In this work, we introduced LHU-Net, a novel light hybrid U-Net architecture tailored for volumetric medical image segmentation. LHU-Net exemplifies the synergy between minimalistic design and advanced performance, utilizing a strategic implementation of attention mechanisms. By integrating spatial attention for initial broad feature capture and channel attention in deeper layers for meticulous feature refinement, LHU-Net efficiently processes diverse datasets and tackles various segmentation challenges with fewer than 11 million parameters. Through LHU-Net, we highlight the untapped potential of achieving high segmentation accuracy at limited model complexity, marking a significant advancement in the development of accessible yet powerful medical image analysis tools.
7 Acknowledgment
The authors gratefully acknowledge the computational and data resources provided by the Leibniz Supercomputing Centre (www.lrz.de).